OpenStudy (anonymous):

An object is moving along a circular path of radius 286 cm. If the centripetal force on the object is three times its weight, find the period of motion

OpenStudy (anonymous):

k wat is centripetal force with formula?

OpenStudy (anonymous):

\[F=(mv ^{2})\div r\]

OpenStudy (anonymous):

yeah there the velocity gives u the tangential velocity at that point in circle so reda ur question and find v then we will talk abt time period

OpenStudy (anonymous):

v = 29.29?

OpenStudy (anonymous):

the wat is time period?

OpenStudy (anonymous):

im confused :(

OpenStudy (anonymous):

61.35?

OpenStudy (mos1635):

F=(mv2)÷r v=2πr/T F=3B=3mg

OpenStudy (anonymous):

time = 6.13 secs?

OpenStudy (mos1635):

g=???

OpenStudy (anonymous):

what g?

OpenStudy (mos1635):

gravitonial acceleration

OpenStudy (anonymous):

go step by step relish the feeling u get when u make ur own discovery of the answer

OpenStudy (anonymous):

why need to solve for gravitational acceleration?

OpenStudy (anonymous):

by using this, F=(mv2)÷r \[3m=(mv ^{2}) \div 2.86m\] v=2.929 m/s then T=2πr/v T=(2π2.86m)/2.929 m/s T = 6.13 secs after this?

OpenStudy (anonymous):

whats gravitational acceleration?

OpenStudy (anonymous):

u have done it that all u got it

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

no problem in the future try following this.....write down the data look for wat u want there will be a slight hint to get what u want to find..... all the best!

OpenStudy (anonymous):

i have another problem, how can i post it?

OpenStudy (anonymous):

did u got it cuty?

OpenStudy (mos1635):

if centripetal force on the object is three times its weight that meens F=3*W W=m*g so that: 3m=(mv2)÷2.86m is not correct

OpenStudy (mos1635):

F=m*V^2/R 3mg=mV^2/R ( bye bye m) 3g=V^2/R 3gR=V^2 V=sqr(3gR) 2πR/T=sqr(3gR) T=2πR/sqr(3gR) if g=9.8 m/sec^2 T=1.958=1.96 sec if g=10 m/sec^2 T=1.939=1.94 sec

OpenStudy (anonymous):

3m=(mv2)÷2.86m is not correct i thought mass = weight .. my bad

OpenStudy (anonymous):

so the correct answer is 1.94 secs not 6.13 secs .. ?

OpenStudy (anonymous):

whats the correct answer here ..